Project

General

Profile

Actions

The MOM Handover List of "Things to Know"

Below is the information that is recorded by the MOM and it is meant to be given to the next MOM.

Paolo Franchini handover to Melissa Uchida (27th February 2017)

Summary of the data taking:
  • TOF calibration
  • Detectors alignment
  • Magnets alignment
  • Momentum scan with 2015-05-1-SSUSSD (FC off)
  • Emittance reduction with 2015-05-1 (different diffuser settings)
Detectors:
  • EMR board replaced with a spare one
C&M:
  • State machine in static mode
Open actions:
  • Spare for EMR (send the old one to Geneva?)
  • Decay Solenoid ALH is disabled
  • Iris 4 is stuck closed in the beamline
  • He in the trackers

Paolo Franchini handover to Paul Kyberd (12th July 2016)

Summary of the data taking:
  • Mock Run MDR7
  • TOF calibration
  • Detectors alignment
ISIS:
  • Beam bump done (28/06/2016)
  • 1.8-2 beam loss
Beamline:
  • Replaced PS of Q1
  • Replaced PS of Q2 after a water dripping on Q123
  • D1 tripped for a MPS Waterflow failure
  • TOF2 channel (TOF_103) at low voltage, tripping at nominal current
  • Downgrade of the DAQ firmware
  • EMR-DAQ problem disappeared after re-cabling the problematic board.
  • 24 V PSU replaced in RR2
Spectrometers/Trackers:
  • Check the Helium flushing in the trackers and in the SSs.
  • Calibration issue during the MDR
FC/LH2:
  • not being used
Control Room Status:
  • Beamline/cooling channel tags
Open actions:
  • TOF2 PMT (Maurizio)
  • Problem in the DAQ (Yordan)
  • ISIS is not reporting beam loss variables to MICE (Pierrick/ISIS)
  • RATS is not reporting Target Depth and Target Delay to the CDB, the fields have a 0.0 (Ed/Joe/Matt)
  • Add monitor in the AppLauncher for the TKU/TKD pressures (Pierrick)
  • Configure the firmware in the spare board of the EMR DAQ (Francois)
  • Hall probe HP-67 is not working (Melissa) - Turnaround in the DAQ to be cleared (Yordan)
  • Fix the "save tag implementation" in RunControl (Pierrick)
  • Physics devil (Scott)
  • Trips of conventional magnets should be reported in the ALH (Pierrick)
  • Trackers helium: can a bottle run for more than 3 days? (John, Craig)
  • More on: Step 4 Readiness (summer 2016) (Steve/MOM)

Paolo Franchini legacy (23rd December)

Summary of the data taking:
  • TOF0/TOF1 alignment run: 1 day
  • FCU @ 100 amps (flip mode): 2 days
  • FCU @ -100 amps (flip mode): 1 day
  • Xenon in the absorber for scattering measurement without field: 4 days
  • Reference run without Xenon in the absorber (replaced by He): 1 day
ISIS:
  • Tune of the beam dump done the 3rd Dec. with Dean Adams
  • 800 Mev, standard dipping was 64/50, several beamloss configurations: from 0.8V to 1.5V
  • Attended the machine Scheduling Meeting of the 16th Dec.; next meeting 26th Jan. 2016
  • ISIS was down nearly 24 hours for a vacuum leak in the Linac
Beamline:
  • Temperature DeltaT readout swap of Q4/Q5 sensors
  • Q6 has still a slight leak; the webcam is useless during data taking (not enough light)
  • Still no DS
Spectrometers/Trackers:
  • He leak in SSD: we had run without He during the scattering measurements
FC/LH2:
  • FC has run in flip mode at 100 amps and -100 amps without any problem
  • Absorber filled with Xenon, evacuated after 6 days and filled with He.
  • SW left. Is VB the new on site expert? She needs an account on MICEmine.
Control Room Status:
  • No major issues that have prevented the normal activities
  • Software versioning procedure used for software updates
  • More in the 'Open actions'
Open actions:
  • Password for experts in the App Launcher #1804
  • Tags in the Run Control for the FCU
  • Shifter version of the Target Drive
  • Instructions to run OnRec needs to be updated on http://micewww.pp.rl.ac.uk/documents/32
  • The Run Control does not write on CDB if a wrong character is inserted in the comment more than two times
  • Pierrick has to send a task list for the network reconfiguration #1803
  • OnMon is not reporting the right status in the Run Status GUI #1802
  • Some spontaneous ends of DATE during the data taking without any error message
  • Everytime the absorber was condidered "empty" was instead filled with Helium?
  • The Physics Devil needs to reach a stable running version
  • Detectors alarm handler has false alarms

Ed Overton to Paolo (2nd December)

ISIS:
  • back at 800MeV. Need to arrange with Dean/Henry to re-tune beam bump [TODO].
  • activation studies will be requested by ISIS
  • Standard dipping is 64/50, 2V.ms beamloss
  • Beam request for December: December 3rd - 17th, 08:00 – 00:00
  • Machine Scheduling Meeting: 16th December, not sure if this is specific to machine physics.
Beamline:
  • Q4 had pipes cleaned. Now operational.
  • Q6 has a slight leak, needs keeping an eye on (webcam).
  • DS Power supply had serious fault and has been craned out. No DS.
Spectrometers/Trackers:
  • Waiting on magnet review.
  • Standing request from pierrick to alert him/record SS quench data in event of an FC quench.
  • Tracker HE leak "repaired", still leaky but probably OK.
    • TOF 1 was removed for this work, has been put back, surveyed and checked with data.
FC/LH2:
  • FC has operated at full current in Flip Mode (225A). Alignment studies await us.
  • LH2 system has two leaks, and will require removal of FC (Estimated 3-6 weeks offline)
Control Room Status:
  • SW Versioning, is being implemented. Need to chase Durga/Pierrick [TODO].
  • Control buttons are easy to click and can cause serious concequences [TODO].
  • Need to chase up Run Control to make sure its ready for data taking.
  • Summary of control room status in MDR here: RunPlan20151124
    • Summary of summary: Most things in good shape. Need to chase Online Monitoring/Reconstruction to improve[TODO]
  • Expert GUI Passwords.
December Running:
  • No LH2 (scattering studies).
  • SSU ECE waiting on US magnet review.
  • Focus Coil Alignment:
    • Need to enure that FC expert is present to drive coil.
    • Need magnetic sweep after closing hall.

Note:

Melissa Uchida to Ed Overton (4th Nov 2015)

DONE
  • Target works successful.
  • FC ramped to 114 Amps in Solenoid mode.
  • Absorber cooled down.
  • Tracker vacuum pipe blew and was replaced on cryo 1. Pumped down successfully.
  • Tracker interlocks moved to RR2
  • Personal oxygen monitors good for 20 months. Press and hold button to start working before use.
  • Tof works on bar 7 compplete.
To DO
  • FC soak text commences wc 2/11/15 [Done EO]
  • SSD leak test
  • Activation run a success, however, activation high and discussion ongoing as to agreed run rate [Done EO]
  • 4V vs 2V run rate TBA with ISIS [Done EO - Now at 2V]
  • Do we need a new activation run as ISIS will now run at 800 [ISIS will monitor and request].
  • Diffuser to be fitted. I have checked that we are good at Oxford and ready to go. Just waiting for a good time to work in the hall. need Antonio present for TOF move. If working after 5 Joe needs 1 week notice to get overtime agreed. Better this way than erecting a tent. Work with Melissa Colin and Geoff and Joe Tacon (Ox) on this. [Done EO]
  • Add veiners and Vesdas to Tracker interlocks in RR2 - Daresbury
NOTE
  • No works to take place at the back of any racks without a risk assessment. Ask DC
  • Meetings recorded in MOM logbook.

Victoria Blackmore to Ryan Bayes (24th June 2015)

  • Almost set up for wireless internet in hall
    • need Andy G. and Craig MacW. to finish installing (2) ethernet cables

*Safety review for magnet powering. Date TBD.

  • ISIS back to 800 MeV at the end of August
    • Activation study to follow during machine physics period.
  • Have calibrated TOFs
    • will need to recalibrate TOFs if moved
  • Tracker alignment data
    • missing muon beams because no DS
    • requested 200 and 300
    • has 270 pi and 190 pi
    • want 200 and 300 MeV muon beams (no particular emittance specified)
    • struggling to get high momentum because of temperature/chiller problem
    • take 70k+ at TOF2 (corrected from 50k due to tracker timing issue)
  • Ask John Govans if he can fix the beam stop hydraulic hose
    • hose is cracked
    • need no beam.
  • Remind John Govans for CM42 CDM training.
  • Update MOM slides on MOM computer (etc.)

*Jaroslaw to be told when the DS is reliably running.

  • Need to follow up with Dean Adams at the start of run period 1b tuning the beam rate.
    • spread out the beam loss over the 3 ms spill gate.
    • Ed must be there!!!
  • Maximum current for quads --- to be informed by Henry's studies.
  • Water pump needs to be pushed independent of magnet cooling (mustn't forget).
    • won't be installed until it is safe to turn the compressors off.
  • Data has not appeared on the grid for select shifts.
    • needs alarm or reminder
  • TOF2 is being noisy
    • test using cosmics or TOF2 trigger.
  • If time during magnet training.
    • check helical work.
    • check magnetic axis.
  • Unlikely if TOF2/KL/EMR to be moved before 10th of July (if they are moved).
  • Need to make sure that all important field off data is taken before they turn fields on.
    • Make sure that Melissa has all of her runs.
  • Determine timing for first quench test.
    • training will commence after that.
  • Ensure simulations of helical tracks are done.
  • The labels on the focus coil turret are reversed check with John Cobb.
  • CDB some values going into run wrong.
    • List to come soon
    • Janusz has a solution for re-writting the CDB but it requires rewriting everything
    • no archival capability.
    • will ask Chris Rogers for code.
    • need to apply pressure for this.
  • TOF1 needs re-surveying
  • there is a survey of focus coil 2
    • if this was from before the magnets were bolted together it will need to be resurveyed.
  • BLOC magnet checklist needs updating with the latest magnet water flows. Remind Henry
  • Remind somebody to check the polarity of the Super-Conducting magnets before they are powered.
*Meetings
  • Daily Ops meetings 6:00pm POD
  • Daily commissioning meetings 3:00pm ISIS coffee lounge
  • Friday run planning meeting: 11:30 am POD
  • beam request valid until Monday Morning (email to Alan Stevens). ATW valid till Monday morning (remind John Govans).

Chris Rogers to Pierrick Hanlet (4th February)

Mock data run (MDR)
  • MDR is written up, summary document is in preparation
Shifter training
  • Check with ISIS that beamline elements are available
  • Check with John Govans how ISIS radiation training should be handled
  • Check that systems are ready to be trained on
  • Send an email reminder
  • Look for the shifter training notes - probably in R76
  • Note that there are a whole load of documentation bits and pieces in RR2
Mock data run 2
  • Experts only run
Practice magnet alignment
  • Chase integration physicist for this
Run 8th March
  • Experts only run
  • Won't be an activation run, but will be data taking
Run 21st/22nd March
  • Steve Boyd to recruit shifters
  • MOM to pursue things are working

Decay solenoid will be going off for maintenance sometime

Documentation review Survey
  • Rogers has two survey notes which he will upload as MICE Notes
  • Think about the rest of the survey and decide how you feel about the situation
  • Consider surveying kit in, depending on PRY status
MOM tasks taken over or delegated
  • Beamline checks during the shutdown goes back down to Henry
Machine Schedule Meeting Tuesday 15:30
Ops meetings every monday morning
  • Craig MacW, Ken Long, MOM, Steve Boyd, trying to get Jaroslaw involved, 10 am Monday

Handover is Wednesday
Operations calendar

Paul Soler to Ryan Bayes (23rd June to 7th July)

  • Decay solenoid power supply
    • FUG will arrive on 23 June (pm) to commission DS power supply
  • Preparations for the 23 June BLM run.
    • Run to take place 9am-12:00pm
    • Test of the BLMs with no beam into the hall (just dipping the target).
    • May involve DAQ test (may be held for activation run).
    • Could not finish locking up of hall on Friday 20 June:
      • Fenced off magnets and fenced off mezzanines (to avoid people moving from one mezzanine to another during search)
      • Crane was not moved towards West end of hall so could not finish securing hall
    • Finish securing hall at 7:30am on 23 June
      • Ensure that problems that occurred in April so not occur again: put person at door to ensure that no unauthorised personnel enters hall while search is ongoing
      • Controlled entry checklist should be followed at all times
  • Preparations for the 29 June run:
    • Run plan uploaded but might need further tweaking to finalise plan

Ryan Bayes to Paul Soler (9th June to 21st June)

  • Locks for the old decay solenoid power supply are not believed to be sufficient for the new incoming power supply
    • A new solution will need to be identified when the new power supply unit will be commissioned in a few weeks.
    • Documentation will need to be updated once the solution is identified.
    • Old decay solenoid PSU control locks and keys should be kept until after new PSU commissioning. The locks are sitting on the desk in the MCLR until a better place is identified.
  • A beamline test will be conducted early in the week of the 9th
    • Dependent on the re-connection of the D2 power supply after the exchange of the DS PSU.
    • Henry Nebrenski will communicate further details.
  • Preparations will be required for the 23 June BLM run.
    • Will only take place on the morning of the 23rd starting at 9am sharp.
    • Will involve a test of the BLMs with no beam into the hall (just dipping the target).
    • May involve DAQ test (may be held for activation run).
    • May require lock up of hall over the weekend of the 21st and 22nd.

Melissa Uchida to Ryan Bayes (12th May to 26th May)

  • Step IV readiness review is ongoing, see Step4RunningRequirements
    • All detector people have been emailed, and also Yordan for DAQ, and had a chat with Pierrick from C&M. This happened two weeks ago and 'in person' at the VC and again 25/5/14. Please email again in a couple of weeks to keep the pressure on. ATM only the tracker have completed this task.
  • BLM and target test day is in the process of being scheduled with Dean Adams.
    • The tests should take about 1 day.
    • Emails have been sent to begin the process from the MOM email.
    • Steve B. has been informed.
    • We need three shifters for the day (2 and 1 to cover breaks) 1 of these can be the MOM. Do we also need a BLOC?
    • Pierrick would like it to be after June 9th when he is back from leave so that the network can be tested. Therefore date suggested has been 9th- 20th June.
  • An issue has been found with the Decay Solenoid PSU CandM window. The DC and AC contactor red boxes do not change state which makes operation confusing. Mike C and Vicky Bayliss raised the issue. Pierrick informed by email but no MICEMine issue raised.
  • There will be no phones available in hall and control room on May 30th. This is a potential danger and must be mitigated
  • The next MICE operations meeting will be Thursday 30th May 2013 at 15:00 BST.

Adam Dobbs to Melissa Uchida (28th April to 12th May)

  • Step IV readiness review is ongoing, see Step4RunningRequirements
    • All detector people have been emailed, and also Yordan for DAQ, and had a chat with Pierrick from C&M
    • Tracker is done, large integration schedule from Craig
    • Lucien and Ruslan have replied saying they will give me something
    • Tracker now the template for what we want
    • Will almost certainly involve arm twisting...
  • Daily walkabout during the week remains Hall team responsibility, if a MOM is staying at RAL then they can do it on weekends instead
  • Oil level checks in Kaeser compressor. See emails in MOM email. MOM to check oil level on weekend as part of walkabout, using webcam. If looks to overflow container then we have a problem.

Adam Dobbs to Ray Gamet (31st March - April 13th)

  • See MOMHandover_2014_04_11.pdf
  • Preparing for Step IV: http://micewww.pp.rl.ac.uk/projects/operations/wiki/Step4RunningRequirements
  • Daily walk about now handled by John Govans - no longer a MOM job
  • DSA opened to allow measurements for new proton absorber system
  • ISIS off until 25th April
  • Work carrying on for Hall AC units (John Govans handling)
  • Moving spectrometer from R9 to Hall expected in last week of April
  • Beam stop juddering repaired on 11-04-14
  • Need to consider how to stop people wandering in to MICE Hall during the hall search (discussed at last hall meeting)

Ryan Bayes to Pierrick Hanlet (October 28 --)

  • Drilling to take place in hall week of 28th of October. Discussion of dust counter-measures to be arranged by phone.
  • Organize/update training records of collaboration members into a useful form.
  • Machine scheduling meeting: 13:00, 6th November, CR03, R61 -- conflicts with collaboration meeting

Ian Taylor to Ryan Bayes (September 30th -- October 30th)

  • MICE-ISIS Safety Meeting: 10:00, 29th October.
  • Expecting .xit files from Tim Hayler, for the survey of D2 in the DSA, which will need to be stored on the MOM laptop. Now a MICE note.
  • Additional training for shifters: All shifters should be told to:
    • get a RAL site pass (or a temporary one from ISIS),
    • be given privileges to enter R5.4 (from ISIS),
    • be shown the location of the Target control rack in R5.4, and
    • be given instructions on what to do should the target fail and the frame start raising during running.
      1. Contact the BLOC, explain that the frame is slowly raising out of the beam. This will have tripped off ISIS.
      2. On instructions from the BLOC, and only when the frame is fully raised, go to R5.4, remove the brass key.
      3. Return the key to ISIS, informing them that they can restart running, all MICE interlocks should have been cleared.
      4. No other steps are to be taken until a BLOC or target expert arrives.

Victoria Blackmore to Ian Taylor (September 2nd -- October 2nd)

  • EMR run plan received from Yordan/Ruslan, requires negotiation with RF installation. Potential for running overnight with RF installation under Controlled Access (?) Requires shifters + BLOC/SOC to be notified, plus submitting run request to Alan Stevens when times have been finalised.
  • FC: New diagnostics should arrive on the 2nd/3rd
  • FC: Next training run on the 5th/6th depending on the arrival and installation of the new diagnostics and the status of the LHe level.
  • BLOC training: 2 hour period is being arranged on Sept 15th for training new BLOCs with the target. Ed Overton will train people on the target, other BLOC training by Henry Nebrensky (time to be arranged with Dean Adams, between 0830 and 1630)
  • Target: Needs a monthly exercise session, 500 pulses due on Sept 18th
  • DSA: Will be searched and locked on the 3rd/4th
  • DSA: DS has been opened, inspected and closed up again
  • DSA: D2, proton absorber and GVA1 have been put back in the beamline, proton absorber is ~3mm upstream of its previous location to avoid catching on the DS
  • DSA: 54mm proton absorber cable guide is now restrained by multiple cable ties so that it raises the absorber fully. The cable ties are plastic and may stretch over time.
  • DSA: D2 and proton absorber need re-surveying before the DS is fixed, otherwise their locations will be unknown for the EMR run

Ray Gamet to Victoria Blackmore (August 20 - September 04)

  • Ed Overton and Paul Hodgson will be at RAL all day Thursday 22nd and, if necessary, Friday 23rd to thoroughly check out target.
  • Geoff Barber will be at RAL on Thursday 22nd, working on tracker. Confirm with him arrangements to re-install proton absorber at end of next week.
  • Have e-mailed Yordan/Ruslan/Alain asking for a detailed run plan for EMR run in October and also details of what statistics they require.
  • DS insulating vacuum was poisoned with ~10mbar He last Friday so should now be up to room temp. Vacuum needs letting up to atmospheric pressure with dry nitrogen. Mike Courthold will begin investigations tomorrow (Tuesday 20th).

Victoria Blackmore to Ray Gamet (August 05 - August 21st)

  • 15 and 54mm proton absorbers are stuck (15mm only because we can't move the 54mm), and the cable is broken to the 54mm absorber.
  • ISIS will contact the MOM if they access the synchrotron before the shutdown.
    • Mike Courthold has asked to be informed in this event so they can do some reconnaissance of the Decay Solenoid prior to the shutdown.
  • FC will be trained again during the week of the 5th (date to be confirmed)
  • ISIS machine scheduling meeting on August 7th at 2pm in conference room 11, R3
  • (Sheffield) target group will visit during ISIS shutdown to test/replace fibres on the target.
  • John Palin visiting to survey TOF2 on August 13th (survey requests must go via Tim Hayler)
  • MOM and Andy Nichols need to collaborate and create a milestone list for the EMR installation and run in October for the next MICO meeting (14th August)
  • TIARA equipment should arrive on August 21st

Chris Rogers to Victoria Blackmore (June 26 - July 25)

Ryan Bayes to Paul Soler (June 1 - June 11)

  • Decay Solenoid chiller will be fixed by Vicky Bayless when the requisite part (temperature sensor) arrives.
    ->Part has arrived needs installation
  • Still need to exercise the DAQ. Will require help from Yordan.
  • Scheduling for training shifts in progress.

Linda Coney to Ryan Bayes (May 2 - May 31)

  • Decay Solenoid chiller is broken - will need to be fixed in order for power supply tests to be completed. Mike Courthold and Vicky Bayless are working with the Daresbury team on the PS.
  • Recent DAQ tests have shown problems - no triggers were recorded. This needs to be debugged with help from Yordan Karadzhov, Chris Rogers, Ian Taylor, and Craig Macwaters.
  • Run Control (C&M) should also be tested. This is best done with help from Pierrick Hanlet(remotely) and Ian Taylor(locally).
  • Tim Hayler is coordinating a review of all survey work in MICE. He needs assistance defining what surveys are needed, what points should be used on each detector, how often surveys are needed, etc.
  • The plan for shifter training needs to be put into action with Paul Kyberd.
  • Target testing in R78 is now active and the system should be checked. In particular, keep an eye on the vacuum as they want to make sure the pump is not failing.
  • Contact Lucien Cremaldi to find out accurate limits for CKOV temperature and humidity.
  • See Web Whiteboard for updated info on Hall conditions.

Ian Taylor to ??? (April 3 - April 20)

  • There will be testing of the R78 power systems, April 27th/28th. This may affect the target testing, however that is currently switched off. Andy Nichols has notified Paul Hodgson.
  • Reminder from Andy Nichols, next quarterly session will be April 23rd 10:30 in G-09.
  • Henry Nebrensky has insisted that the Roof Conductivity check be re-instated. He has provided e-mail documentation that there is no undue risk associated in going up there, but hopes that MOM's will be careful, stick to the red walkway and watch for trip hazards.
  • Web Whiteboard. In a push to improve change/config control, important hardware or software configurations will be marked on the 'Web Whiteboard'. Examples would include: "archiving of decay solenoid data disabled during warm period" or "DS interlock bypassed during work". Pierrick Hanlet and Mike Courthold are working on a persistent layout and current draft.
  • Item added to MOM walkabout: [camhall 17] Linde Fridge Compressor shows an oil bottle, which should be mostly empty. In the past, this has suddenly filled rapidly over a couple of days. Monitor bottle, and inform Mike Courthold or Vicky Bayliss if filling rapidly or past 50% point.
  • Ed Overton will probably come down the week of the 22nd, to exercise the target while ISIS is still open.

Adam Dobbs to Andy Nicholls / Pierrick Hanlet (Feb 18 - March 20)

  • ISIS do not wish us to access the MICE hall roof routinely due to safety concerns (falls etc, not radiation). Therefore the magnet water conductivity reading has been removed from the MOM daily walkabout list until / if a remote readout system can be found (Andy looking into it).
  • miceonrec02 is accessible to Matt Robinson. onrec03 still needs to be connected and setup.
  • An extra check has been added for R9. The server micecss2 in the controls rack should be checked to make sure it is running. MOM hand-off checklist and walkabout sheet updated accordingly (in MOM folder in R76).
  • We are still required to check the heater in Lab 7 is running on Saturdays and Sundays.
  • Power adapter for live beam loss monitor in MLCR needs to be replaced. CMacW looking in to it.
  • The OKI printer in R76 is being retired. Its replacement is a new HP colour laser jet which is already setup in R76.

Yordan Karadhov to Adam Dobbs (Feb 5 - March 20)

  • Finish the installation of miceonrec2 and miceonrec3 in MLCR. When this is done contact Matt Robinson and ask him to install Linux.

David Adey to Yordan Karadhov (Jan 7 - Feb 6, 2013)

  • Monitoring of CKOV HV reads last on values despite CAEN crate being turned off
  • Lights in the fire exit over the ISIS Linac still do not work. Also watch for water leaks.
  • Yearly PPS functional tests (C Macwaters aware)
  • ISIS beam loss display in MLCR fuzzy
  • Find out if it's possible to get the cooling water conductivity without going onto the roof (H Nebrensky aware, needs to speak with J Govans)
  • NIM crates in rack room occasionally turning themselves off
  • New survey needed after CKOV removal and replacement in January

Yağmur Torun to David Adey (Dec 3-23, 2012)

  • TOF1 frame interferes with installation of upstream SS support
    • asked Andy not to move it until it's confirmed to have well-understood survey data in its current position; if not, it should be surveyed again
  • Keep an eye on water dripping from ceiling in fire exit hallway
  • Attend ISIS machine scheduling meeting Wed Jan 30, 14:30 GMT, CR16 (R80)
    • make sure MICE activation study (Feb 13) doesn't fall off the list
  • Recruit shift crew for activation study (including BLOC, SOC, DOC?)
  • Update expert list
    • Ruslan should replace Cadoux for EMR (get info from Alain)
    • Stewart no longer hall manager (Andy filling in until replacement)
  • Shifter training
    • Remind Craig M to sign off on PPS part of the forms (for Paul S and YT) in MLCR shifter training folder
    • Forward information about newly minted shifters to Linda or update list on wiki page
    • Consider scheduling a training session a day or 2 before activation study; may convince a few people to arrive early for cm35 and take the training
  • Broken controls
    • TOF/KL GUI has internally inconsistent display about ON/OFF channels; check manually before next run
    • Neutron monitor readout is working but requires update to machine configuration to survive across reboots (Ian Taylor to implement 1st week of Jan)
  • DAQ is broken
    • Schedule testing with Yordan K long before next run
    • Run Control hook to DAQ was disabled because it was triggering a latent bug in DATE (4k spill limit); requires decision by online group
  • DSA access
    • Lucien and David will be at RAL Jan 23-27 and would like CKOV-a to be removed on the 23rd or 24th and reinstalled Jan 24 or 25
    • Coordinate with Andy to make sure surveys are done in the right order (TOF0 and CKOV before/after removal, etc.) and set priorities; include TOF1 if needed
    • Watch out for compressor removal near DSA entrance which may interfere with access
  • PPS readout
    • Ask Craig M to have crate 10 reset to get the mimic working correctly again
  • Hall A/C units
    • Andy is working with John Govans to have the refrigerant lines with bad joints repaired. This needs to be finished before DS operation so we can turn on the units without fear of another leak.
    • Pierrick was supposed to talk to John G who claimed remote status readout was interfering with control of the units. Check to see if this has been resolved.

Henry Nebrensky (Nov. 2012)

  • moved from MOM laptop to MOM federal (roaming) profile
    • Official MOM e-Mail client is Thunderbird
    • Official MOM web browser is Internet Explorer
    • possible issues with details of printers etc.
      • MICE_R1 is missing, CutePDF is broken, need easy way to reach on-line calendar, does it work on other PCs?
    • check what's missing that I didn't think of, and make a list somewhere for the next time ( #1217 )
  • chase Hall A/C units - 3 is down and 1 has an error, and circuit for 4 had a Freon leak
  • Hall safety
    • no lights in fire escape
    • possible asbestos tiles in trench
  • RA/MS - the system still needs setting up. Meanwhile
    • "boilerplate" for those associated with running are on the IIT site
    • print out as needed and put in folder with ATW and EPTW
    • amend with pen if needed. This paper copy is the one that's actually in force.
  • don't let Dell off the hook regarding miceecserv warranty quote
  • schedule time with RAL Networks to install switch - avoid clash with AFC cooldown
  • water seepage into trench - Andy Scott to check it's not from Linac
  • Look into why text messages didn't send from MOM phone after Freon leak

MOM: Ray Gamet (8 Oct to 6 Nov 2012) handing over to Henry Nebrensky

  • Work through MOM-assigned Operations issues in micemine.
  • ISIS
    • respond to ISIS schedule meeting email – note when meetings will happen
    • attend MICE-ISIS Safety Management meeting on 26/11
  • Data Taking
    • Check online reco working OK with cosmic run - provisionally scheduled with Chris Rogers for Friday 9/11
    • If all OK decide when/if to have data-taking run to complete proton absorber studies, recruit shifters/bloc etc.
  • Magnetic Field Group
    • if needed, work with Luke Fry on list of sensitive equipment for Magnetic Field group
  • Safety
    • new RA/MS produced for HV, BS, conventional magnets. Assess need for any others
    • develop system for RA/MS in conjunction with Project Manager/Hall Manager/head of Operations
    • work with Online Group to reinstate remote readout of Neutron Monitor if possible
  • FC Commissioning
    • assist (when needed) with FC commissioning
  • Beamline
    • Ensure appropriate AC units running during DS testing - Mike Courthold will see to this.

MOM: Linda Coney (26 September to 10 October 2012)passed on to Ray Gamet

  • Preparation for Running 12 - 14 October
    • Read MOM instructions – MOM 2012 binder and safety information binder
    • Print out MOM checklist (on MOM laptop) for running
      • Review with Linda
    • Print out target checklist (on MOM laptop) – give to BLOC
    • Work through Hall Preparation list from micemine wiki
    • Give Ian & Yordan shifter training presentation – slides are on MOM laptop
      • Must be done before practical training starts (ie. before noon on Friday)
    • Add updated PPS docs to binders in MLCR
    • Add updated moveFiles docs to binders in MLCR
    • Add updated Beam Stop docs to binders in MLCR
    • Write RA/MS for luminosity monitor
    • Write RA/MS for HV operation
    • Find out status of Neutron Monitor
    • Get Authorization to Work from Andy
      • Note: RA/MS are pre-requisite for this
    • Post appropriate signs on equipment and/or doors in Hall
    • Do test of target operation with BLOC
    • Find latest ISIS schedule from ISIS website
  • Work through MOM-assigned Operations issues in micemine with the help of Operations Head (http://micewww.pp.rl.ac.uk/projects/operations/issues)
  • ISIS
    • respond to ISIS schedule meeting email – note when meeting will happen
    • decide if we want/could to run Activation Run during next Machine Physics
  • Magnetic Field Group
    • work with Luke Fry on list of sensitive equipment for Magnetic Field group
    • take background measurements of field in Hall
  • Safety
    • assess need for other RA/MS
    • develop system for RA/MS in conjunction with Project Manager/Hall Manager/head of Operations
    • work with Online Group to reinstate remote readout of Neutron Monitor
  • FC Commissioning
    • assist in preparation of R9 for FC arrival
    • assist (when needed) with FC commissioning
  • MOM walkthrough
    • clarify purpose of air receiver in Trench
    • is it necessary to check daily? - No, but needs checking weekly so leave on walkthrough
    • Find out when Target testing restarts – add to walkthrough
    • During DS testing – add DS fridge items back to MOM walkthrough
    • Work toward reducing items actually in Hall on MOM walkthrough – get help from Online Group
  • Beamline
    • Schedule Decay Solenoid fridge testing – assist when necessary
    • Ensure appropriate AC units running during DS testing
  • Online
    • learn from Pierrick about new Monitoring applications

MOM: Henry Nebrensky (3 January to 3 February 2012)

  • check with Hall Manager that the RA/MS paperwork covering powering of Q789 is appropriate - wide variation in how people treat this. Note crude measurements in eLog 1961 / 1977.
  • review key handling arrangements with Hall Manager and with Operations Group, #895.
  • chase target webcam permissions #894.
  • track neutron monitor discussion #884.
  • monitor parking situation outside Hall - can we get a reserved space for the MICE van?
  • shut down HV supplies (SY 527/127) if necessary - need to understand plumbing work and movement of A/C unit 5. Liase with Hall Manager and supply owners.
  • decide on future of big OKI printer - buy more toner? Liaise with Craig and Matt Robinson
  • On 19th Feb starting midnight, ISIS will run an activation study on Sector 7 using dosimeter badges - no MICE target, this sets background level. Be aware that they'll want a follow-up run with the target dipping: check with Ken, would be either end March or start May after shutdown.
  • We are approaching the regime where we have strong magnets in the Hall and must control movements of magnetic material in and out - need to push DL regarding diagnosis and repair of the metal detector. How do we know only the PSU is damaged? And once it's working again, who is responsible for defining / setting / checking the correct operating parameters?

MOM: Alain Blondel and Victoria Blackmore (14 January to 15 February 2011)

  • follow up on PPS and rebuild of MICE control room power supplies
  • get from ISIS exact dates of running periods for 2011 including the MPs and running up periods
  • follow up schedule of TOF1 reconstruction, EMR and activities to be foreseen in the MICE hall in 2011.
  • by 15 February a program of work in the MICE beamline/hall should be completed for proposal to the collaboration

MOM: Pavel Snopok (3 Oct. 2010 till 9 Nov. 2010)

MOM: Linda R. Coney (10 Sept. 2010 till 3 Oct. 2010)

  • Keys to Q1, Q2, Q3, D1 power supplies are UP IN ISIS CONTROL ROOM. ACCESS DETERMINED BY ISIS DURING SHUTDOWN.
  • Measure polarity on D2, Q4-6, Q7-9
  • Move power strip sitting on floor behind Q123 power supply to avoid potential damage if PS cooling water leaks.
  • Document new van procedures.
  • Bug DAQ expert to update documentation.

MOM: Jean-Sebastien Graulich (8 Sept. 2010 till 11 Sept. 2010)

  • TOF0 and TOF1 were surveyed. Measurements were taken to determine tilt(s) of TOF1 and TOF0 wrt the beamline. MICE Note describing detector positioning and requirements for reinstallation is needed.

MOM: Roumen Tsenov (13 Aug. 2010 till 8 Sept. 2010)

  • Check-list for MOM's Daily-walk-through during the shut-down: I've been using the run-state one, but in fact many items are not needed.
    Mike Courthold should give his recommendations for the cryogenics. Press on him! Updating of the rest is obvious, more or less. Should be done after Mike's updating.
    File (Rev1) in Operaions/CheckLists
  • Compile note on detector geometry.
    Main issues are the positions of TOF0/1/2, KL and GVA1 during the user's run July-Augist 2010.
    Measurements with Faro Laser Tracker are planned for 11-12 Sept. 2010. Talk to Willie.
    Results have to be confronted with what is in the geometry files in the software (Chris Rodergs).
    Special directory in MOM's laptop MICE-OM/Operations/Geometry. Have a look there.
  • (Long range) Documentation must be updated well before ISIS start in March 2011. Many of the manuals contain not up-to-date info, or they are incomplete. Insist on very detailed description. Manuals should be written in a way that an intelligent person/physicist knowing almost nothing about MICE can understand what is where and what to do.
  • MOM's mail folders. Old READ and SENT messages moved to Local folders in MOM's laptop. Only relatively recent READ and SENT are kept on the server. Idea behind that: keep the MOM mailbox size below 1 GB.
  • MOM's laptop Desktop cleaned. Files moved to respective folders according to their content.
  • Desktop on the windows machine in MLCR cleaned, too. All files there moved in a folder "CollectedFromMLCR Desktop16082010" that is on the Desktop. Also copied to MOM's Desktop.
  • New laser printer/scanner/copier Samsung SCX-4600 available in R76 for MICE. Installed on MOM's laptop.
  • Repaired PSU for Q3 installed back by Martin Hughes and Willie's crew. Not tested. Must be done when no work in ISIS vault is going on.
  • Ash things -clear the space

Updated by Franchini, Paolo almost 6 years ago · 166 revisions